6 g of dolid magnesium sulphate, MgSO4 is added into a beaker containing 200 dm3 of water. calculate the concentration in g dm-3, for the solution produced


1
Expert's answer
2021-07-09T05:16:04-0400

Molar Mass of MgSO4 = 120.366

= 6/120 = 0.05M

Concentration = m/Dm3

= 0.05/200

= 0.00025mol/dm3

or

0.03g/dm3